A man wants to make up 100 one-pound jars of nuts to sell for 66 cents a jar. He plans to use 40 pounds of nuts worth 75 cents a pound. What should be the price of the other type of nuts he uses?
Answer by ankor@dixie-net.com(22740)

A man wants to make up 100 one-pound jars of nuts to sell for 66 cents a jar.
He plans to use 40 pounds of nuts worth 75 cents a pound.
What should be the price of the other type of nuts he uses?
:
Since we already know we have 40lb, and we want 100lb, 60lb of unknown price nuts are required
:
let p = the price per pound of 60lb of nuts
:
The mixture equation
75(40) + 60p = 66(100)
3000 + 60p = 6600
60p = 6600 - 3000
60p = 3600
p = 360/60
p = 60 cents a lb nuts required
:
:
:
Check this
75(40) + 60(60) = 66(100)
3000 + 3600 = 6600